Session: Nov 1997 (session 4)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Zacchiroli at Fri Nov 14 15:45:35 MET 1997) Success

Onsala (A. Polatidis at Fri Nov 14 17:08:59 MET 1997) Success
No problems known

Westerbork (tony foley at Sat Nov 15 15:31:31 MET 1997) Minor failures
No phasecal. No other known problems

Jodrell (Alastair Gunn at Sun Nov 16 15:09:57 MET 1997) Minor failures
FS crash at 09:37, restarted recording at 10:36. Track 21 dead. No other known problems.

Torun (K.Borkowski at Sun Nov 16 19:17:39 MET 1997) Minor failures
At 7:14 UTC there was a power failure. Afrer the restart of obserwations some optional settings were incorrect for these observations but remained unnoticed by the operator, eg. wrong frequency (18 cm instead of 6 cm) until 9:47; at 10:26:50 the formatter time was corrected by about 1 s. At 11:19 for about 10 minutes there were problems with the telescope control.

Noto (C.Trigilio at Tue Nov 18 13:37:33 MET 1997) Minor failures
Lost some minut for cable wrap problem at 6:50.

Effelsberg (dgraham at Thu Nov 20 15:40:46 MET 1997) Success

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:04:34 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name v014b1.antab Read the comments in the file to get further information.